OpenStudy (anonymous):

What 2 numbers are the product of -12 and the sum of -23

hero (hero):

No such number exists.

hero (hero):

Actually, it does exist, but the numbers are not integers

OpenStudy (anonymous):

the are integers they are just odd

hero (hero):

Negative.

OpenStudy (anonymous):

yea srry lol

hero (hero):

\(-11.5 + \frac{\sqrt{577}}{2}\) and \(-11.5 - \frac{\sqrt{577}}{2}\)

hero (hero):

Those are the numbers

OpenStudy (anonymous):

oh wow how did you get that

hero (hero):

Magic methods

hero (hero):

You want the formula?

OpenStudy (anonymous):

yes please!

hero (hero):

Well, I can't just give it to you, I have to explain it to you.

OpenStudy (anonymous):

of course.

hero (hero):

You're working with trinomials or quadratic equations, right?

OpenStudy (anonymous):

yeah

hero (hero):

I'm trying to figure out why you would be given such a problem. Is this the entire problem you were given? (Probably not)

OpenStudy (anonymous):

yeah it is. and I'm not sure why either. :/

hero (hero):

Weird, but okay. So for an explanation, I'm going to use vyew.

OpenStudy (anonymous):

wait! my bad. it isn't the whole problem.

OpenStudy (anonymous):

sorry, the whole problem is:

hero (hero):

lol, I told you :P

OpenStudy (anonymous):

Factor 24x^2 - 23x - 12 I have to find the product of the last number and the sum of middle number.. I did exactly that... but I couldn't find any number that fit the two. :/

hero (hero):

I can show you in vyew how to do it. Do you have headphones?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

is vyew on this website?

hero (hero):

Okay, hang on a sec

hero (hero):

No, it's another website that has whiteboard and voice capability

OpenStudy (anonymous):

okay can you send me the link to the website or website name please?

hero (hero):

I will send you a link. Hang on

OpenStudy (anonymous):

okay